PITMinutes 09-04-2018

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, K. Baggett, E. Daly. A. Freyberger, R. Geng, J. Guo, C. Hovater, A. Hutton, G. Krafft, A. Kimber, R. Legg, F. Marhauser, C. Mounts, R. Nelson, T. Plawski, T. Powers, C. Reece, T. Reily, R. Rimmer, and A. Seryi.

Agenda and Notes

1. Operations Status and Summary/Review (Freyberger)

Multipass beam reestablished in preparation for the physics run. In the middle of last week beam to Hall C and CW checks were initiated. Beam energy of 1050/linac was achieved with fairly reasonable trip rates as the RF recoved (see 2.)

Toward the end of last week vacuum issues, especially in the north linac SRF, necessitated halting of beam delivery to physics. Investigations of leaks continuing.

1L21 had condensed gasses but no clear indications of external leak. Stay tuned.

RF Heat and Q0 are probably not in balance. Indicated by different JT position RF on/off. Start balancing out during the week. Will do all 10 C100 modules. SLINAC at 2K means can start south side this week.

2. CEBAF Turnon Status (Solopova)

C100 Status Before Vacuum Troubles

3. Recent Microphonics Measurements (Powers)

C100 System Improvement Discussion

Items Request Email

C100 Refurbishment Brainstorming List

Tony's C100 Refurbishment Brainstorming ppt

4. Beam Pulsing for JLEIC (Guo)

Beam Pulsing Measurements at CEBAF

Action Items

Leadership of quenches. Offline discussion.

C100 Refurbishment Project for the future?